Comprehensive CVP analysis (LO 1, 2.3, 5) Tll never understand this accounting stuff. Blake Dunn yelled, waving the income statement he had just received from his accountant in the morning mail. Last month, we sold 1,000 stuffed State University mascots and carned $6,850 in operating income. This month, when we sold 1.500, I thought wed make $10275 But this income statement shows an operating income of $12,100! How can I ever make plans if I cant predict my income? Im going to give Janice one last chance to explain this to me, he declared as he picked up the phone to call Janice Miller, his accountant Will you try to explain this operating income thing to me one more time? Blake asked Janice. After I saw last months income statement, 1 thought each mascot we sold generated $6.85 in net income; now this month, each onc generates $8.07! There was no change in the price we paid for each masoot, so I dont understand how this happened. If I had known I was going to have $12,100 in operating income, I would have looked more seriouslyat adding to our product line Taking a deep breath, Janice replied, Sure, Blake. Id be happy to explain how you made so much more operating income than you were expecting. Required Assume Janices role. Explain to Blake why his use of operating income per mascot was in error b. Using the following income statements, prepare a contribution margin income statement for March February March Sales revenue Cost of goods sold Gross profit Rent expense Wages expense Shipping expense Utilities expense Advertising expense Insurance expense Operating income $25,000537,500 10,000 15,000 15,000 22.500 1,500 1,500 3,500 5,000 1,250 1,875 750 750 750 875 400 400 6,850S12,100 Blake plans to sell 500 stuffed mascots next month. How much operating income can Blake expect to eam next month if he realizes his planned sales? d. Blake wasnt happy with the projected income statement you showed him for a sales level of 500 uffed mascots. He wants to know how many stuffed mascots he will need to sell to eam $3,700 in operating income. As a safety net, he also wants to know how many stuffed mascots he will need to sell to break even. Blake is evaluating two options to increase the number of mascots sold next month. First, he believes he ean increase sales by advertising in the university newspaper. Blake can purchase a package of 12 ads over the next moeth for a total of $1,200. He believes the ads will increase the number of stuffed mascots sold from 500 to 960 A second option would be to reduce the selling price. Blake believes a 10%% decrease in the price will result in 1,000 mascots sold. Which plan should Blake implement? At what level of sales would he be indifferent between the two plans? Just after Blake completed an income projection for 1,200 stuffed mascots, his supplier called to inform him of a 20% increase in cost of goods sold, effective immediately. Blake knows that he cannot pass the entire increase on to his customers, but thinks he can pass on half of the 20% increase while suffering only a 5% decrease in units sold. Should Blake respond to the increase in cost of goods sold with an increase in price? g. Refer back to the original information. Blake has decided to add stadium blankets to his product line. He has found a supplier who will provide the blankets for $32, and he plans to sell them for $33. All other variable costs currently incurred for selling mascots will be incurred for selling blankets at the same rate. Additional fixed costs of $350 per month will be incurred. He believes he can sell one blanket for blankets and stuffed mascots will Blake need to sell each moeth in order to break even? every three stuffed mascots. How many
Answer #1

a. Janice should explain Blake in following manner :-

" Blake your are making a technical error in understanding the concept of operating income . Operating income can be calculated on a per unit basis exactly only when all the costs remain variable. But in case any of the cost is fixed then it will definitely result in a different operating income figure. Sales minus all variable cost is known as contribution. And as and when contribution exceeds fixed cost we start making profit and that's why operating profit on sale of 1500 units is coming out to be different than what it was at sale of 1000 units.

(b) . Contribution Margin income statement will be as follows :-

Sales revenue 37500
Less .:- variable expense
- Cost of goods sold (15000)
- Wages (1500x3) [note 1] (4500)
- Shipping exp (1875)
- Advertising exp (0.25x1500) [note 2] (375)
Contribution $15750
Less : fixed cost
- wages exp [note 1] (500)
- advertising exp [note 2] (500)
- rent exp (1500)
- utilities exp (750)
- insurance exp (400)
Operating income $12100

Working notes

1 . Variable portion of wages expense

= Change in cost / change in units

= 5000-3500/1500-1000

= 3 per unit

Thus fixed portion of wages exp = 5000-3x1500 = 500

2 . Variable portion of advertising exp = 875-750/1500-1000 = $0.25 per unit

Fixed portion = 875-0.25x1500 = $500

C. Contribution per unit = total contribution / units sold

= 15750/1500 = $10.50 per unit

Fixed cost = $3650.

Operating income at the sale of 500 units = 10.50x500-3650 = $1600

At the level of 500 units operating profit will be $1600.

D . Total contribution if required operating profit is $3700 = 3700+3650 = $7350

Toal units required to be sold ato achieve desired level of contribution = desired level of contribution / contribution per unit

= 7350/10.50 = 700 units

For breakeven required number of stuffed mascots to be sold = fixed cost / contribution per unit

= 3650/10.50 = 347.62 or 348 units

Note :- due to lack of time on board , I am able to solve this much, however I tried to solve what is minimum required from us i.e. 4 sub parts. Please upload the remaining parts, I will definitely complete them then. I am extremely sorry for the inconvenience caused to you. I hope you will understand.